Math 550 (Section 1)  University of South Carolina
D. Meade  Fall 1997

Day One Questionnaire – Math 550 (Fall 1997)

• Personal Information Name Phone Number E-mail Address Major Year

• \underline{Schedule}

Work (hrs/wk) Relax (hrs/wk) Study (hrs/wk) Math 550 (hrs/wk) Other Courses

• Mathematical Background

Course  When  Where  Who  Grade

Math 142 (or equiv)

Math 241 (or equiv)

Other Post-Calculus Math